OpenStudy (anonymous):

12(2x - 8) = 4(5 - 2x) + 3(7 - 3x)

OpenStudy (anonymous):

24x-96= 20-8x+21-9x 24x-96=-17x+41 41x=137 x=3.34 maybe

OpenStudy (anonymous):

thanx n how u put 3.34 in fraction?

OpenStudy (anonymous):

3 17/50 which is the same as saying three and seventeen fiftieths

OpenStudy (anonymous):

gracias

OpenStudy (anonymous):

No problem c~: